OpenStudy (nikki_skellington):

The sets of numbers 3, 4, 5 and 8, 15, 17 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

OpenStudy (steve816):

\[\Huge a^2+b^2=c^2\]

OpenStudy (steve816):

Pythagorean theorem ^

OpenStudy (nikki_skellington):

so 3^2+4^2=5^2 is the first one....

OpenStudy (steve816):

Yup! That is a Pythagorean identity because the equation is true! Now, do it for the other ones!

OpenStudy (steve816):



OpenStudy (nikki_skellington):

8^2 +15^2 =17^2 right?

OpenStudy (steve816):

Exactly!

OpenStudy (nikki_skellington):

ohhh ok i get it now, thank you!

OpenStudy (steve816):